OpenStudy (anonymous):

if f(x)=1/x-3 and g(x)=f'(x) find g'(x)

OpenStudy (bibby):

this is asking for the inverse of the inverse. This is probably f(x) logically. Start by finding the inverse of f(x)

OpenStudy (anonymous):

so g'(x) = f'' (x)

OpenStudy (bibby):

I might actually be wrong. Anyhow. \[y = \frac{1}{x-3}\] \[x = \frac{1}{y-3}\] solve for y

OpenStudy (bibby):

inverse or derivative?

OpenStudy (anonymous):

derivative. it is the derivative of the derivative. I got it. thx. I have another question exactly the same except g(x)= f^-1 (x)

OpenStudy (anonymous):

it's asking for 2rd derivative of f

OpenStudy (bibby):

she already got it :)

OpenStudy (anonymous):

so what do we need to find? g'(x)?

OpenStudy (anonymous):

or g(x)?

OpenStudy (anonymous):

if f(x)=1/x-3 and g(x)=f^-1 (x) find g'(x)

OpenStudy (anonymous):

So lets find the inverse and follow what bibby wrote above

OpenStudy (anonymous):

\[y = \frac{1}{x-3}\] Then we switch our x and y \[x = \frac{1}{y-3}\] Then Solve for y \[y=\frac{1}{x}+3\] so \(g(x)=\frac{1}{x}+3\)

OpenStudy (anonymous):

Now find the derivative of g(x)

OpenStudy (anonymous):

\[g'(x)=-\frac{1}{x^2}\]

OpenStudy (bibby):

remember that you can add derivatives so \[g'(x) = \frac{d}{dx}\frac{1}{x} +\frac{d}{dx}3\]

OpenStudy (anonymous):

1/3x^2???

OpenStudy (bibby):

the derivative of 3 (a constant) is 0

OpenStudy (bibby):

you can use either the power rule because 1/x = x^-1 or the division rule

OpenStudy (anonymous):

1/x^2?

OpenStudy (bibby):

-1/(x^2) \[x^n - > nx^{n-1}\] \[\frac{ d }{ dx }x^{-1}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok like what dont u follow?

OpenStudy (bibby):